OpenStudy (anonymous):

find the domain and the range of y=square root of x-5

OpenStudy (anonymous):

For\[x\in [5,\infty]\], \[y\in\mathbb{R}\] and for \[x\in[-\infty,\infty]\], \[y\in \mathbb{C}\]